2 years ago, Greg was 3 times older than his brother. In 4 years, he will only be 2 times older. What are their present ages?
Answer by Paul(988)

3x+x-2=4+2x+x
-2+4x=4+3x
x=6
3(6)=18
Hence, Greg is 18 and his brother is 6.